OpenStudy (anonymous):

If during take off a 40,000 kg aircraft accelerates at 4 m/s^2, what force is being generated by the engines?

OpenStudy (anonymous):

force=ma m=mass a=acceleration

OpenStudy (anonymous):

so whats the answer@cherio

OpenStudy (anonymous):

@cherio12

OpenStudy (anonymous):

40000*4

OpenStudy (anonymous):

ohh okay i had that right can you help me with one more

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

An 80 kg man starts from rest and after 3.50 seconds is running at a speed of 12.0 m/s. The force needed for this man to get up to this speed in 3.50 seconds is ____ N. @cherio12

OpenStudy (anonymous):

vf=vo+at vf=final velocity (12m/s) vo=initial velocity (0m/s) a=acceleration (what you want to find) t+time (3.5s)

OpenStudy (anonymous):

80*a=F

OpenStudy (anonymous):

so the answer is 80*a=F

OpenStudy (anonymous):

or is it 80

OpenStudy (anonymous):

80N

OpenStudy (anonymous):

80*a=F F is the answer you want

OpenStudy (anonymous):

SOO The force needed for this man to get up to this speed in 3.50 seconds is __F__ N

OpenStudy (anonymous):

well yes..but you need to plug in the values

OpenStudy (anonymous):

HOW DO I DO THAT

OpenStudy (anonymous):

I showed you all the values and where they go in an earlier post

OpenStudy (anonymous):

YEA I KNO BUT HOW DO I SOLVE IT DO I MULTIPLY

OpenStudy (anonymous):

12=a*3.5 a*80=F

OpenStudy (anonymous):

a=3.4286

OpenStudy (anonymous):

A=F/80